# Adaptive Coded Federated Learning: Privacy Preservation and Straggler Mitigation

### v1

@@ Abstract
In this article, we address the problem of federated learning in the presence of stragglers. For this problem, a coded federated learning framework has been proposed, where the central server aggregates gradients received from the non-stragglers and gradient computed from a privacy-preservation global coded dataset to mitigate the negative impact of the stragglers. However, when aggregating these gradients, fixed weights are consistently applied across iterations, neglecting the generation process of the global coded dataset and the dynamic nature of the trained model over iterations. This oversight may result in diminished learning performance. To overcome this drawback, we propose a new method named adaptive coded federated learning (ACFL). In ACFL, before the training, each device uploads a coded local dataset with additive noise to the central server to generate a global coded dataset under privacy preservation requirements. During each iteration of the training, the central server aggregates the gradients received from the non-stragglers and the gradient computed from the global coded dataset, where an adaptive policy for varying the aggregation weights is designed. Under this policy, we optimize the performance in terms of privacy and learning, where the learning performance is analyzed through convergence analysis and the privacy performance is characterized via mutual information differential privacy. Finally, we perform simulations to demonstrate the superiority of ACFL compared with the non-adaptive methods.
@@ I Introduction
In intelligent systems, edge devices are generating an increasing amount of data, which can be utilized to train machine learning models for various tasks across a wide range of applications, such as smart cities [1], autonomous driving [2], and intelligent healthcare [3]. Traditionally, these data are uploaded to a central server for centralized machine learning. However, uploading datasets directly to the central server without privacy-preservation measures raises severe privacy concerns for the edge devices [4]. As an alternative to centralized machine learning, federated learning (FL) has emerged as a highly effective tool [5, 6, 7]. Under the FL framework, during each iteration of the training process, local devices compute gradients with their local datasets and send these gradients to a central server. The central server then aggregates the received gradients to update the global model [8]. In this manner, the edge devices only share locally computed gradients with the central server, significantly enhancing the privacy of raw datasets.
@@
In FL, the training process may be hindered by devices known as stragglers, which are significantly slower than others or completely unresponsive due to various incidents [9, 10]. Additionally, in FL, the local datasets on different devices can be highly heterogeneous, meaning they are not independent and identically distributed (non-i.i.d.) across devices [11, 12]. In this case, the presence of stragglers can bias the trained global model due to the absence of the contributions of some devices in each iteration. In other words, stragglers can negatively impact the learning performance of FL, potentially leading to poor convergence. In traditional distributed learning scenarios, the challenges posed by stragglers can be addressed through well-known gradient coding (GC) strategies [13, 14, 15, 16, 17, 18, 19]. The core concept of GC is replicating the training data samples and distributing them redundantly across the devices prior to training. During each training iteration, non-straggler devices encode the locally computed gradients and transmit the encoded vectors to the central server. Upon receiving these vectors, the central server is able to reconstruct the global gradient, effectively mitigating the impact of stragglers. However, in FL scenarios, privacy concerns make it impractical to adopt GC techniques and to allocate training data to devices redundantly in a carefully planned manner, as is done in traditional distributed learning scenarios. This is because devices are reluctant to share their local datasets directly with others to create data redundancy considering that their local datasets may contain sensitive or private information [20].
@@
For FL problems, to evade the negative impact of stragglers, a new paradigm named coded federated learning (CFL) has been investigated [21, 22, 23, 24]. In CFL, each device uploads a coded version of its dataset to the central server to form a global coded dataset before training begins. During the training process, the central server aggregates gradients received from the non-stragglers and gradient computed from the global coded dataset to mitigate the negative impact of the stragglers. Compared with sharing raw datasets as in GC, by carefully designing the way of encoding local datasets in CFL, privacy concerns can be largely reduced, and one can even achieve perfect privacy. However, it is intuitive that encoding the local datasets induces information loss and this may negatively impacts the learning performance. Based on that, the key issue in CFL is how to generate the global coded dataset before training, as well as how to aggregate the gradients at the central server during the training process, in order to achieve better performance in terms of learning and privacy.
@@
There has been some recent progress under the framework of CFL. To be more specific, CFL was firstly investigated in [21] for linear regression problems under the FL setting, where local coded datasets are constructed at the devices by applying random linear projections. In [22], the CFL scheme was combined with distributed kernel embedding to make CFL applicable for nonlinear FL problems. Later, to further enhance the privacy performance of CFL, a differentially private CFL method was proposed by adding noise to the local coded datasets, where the differential privacy guarantee was proved [23]. Very recently, a stochastic coded federated learning (SCFL) method has been proposed [24]. In SCFL, before the training starts, the devices generate and upload privacy-preservation coded datasets to the central server by adding noise to the projected local datasets, which are used by the central server to generate the global coded dataset. In each iteration of the training process, the central server aggregates gradients received from the non-stragglers and gradient computed from the global coded dataset with fixed aggregation weights. Compared with previous methods, the advantages of SCFL lie in ensuring that the global gradient applied by the central server to update the global model is an unbiased estimation of the true gradient, and providing the trade off between the privacy performance and the learning performance analytically. However, in SCFL, when the central server aggregates the gradients, it consistently adopts pre-specified and fixed weights across iterations, which neglects the generation process of the global coded dataset and the dynamic nature of the trained model over iterations. This oversight results in an unnecessary performance loss, preventing SCFL from achieving the optimal performance in terms of privacy and learning.
@@
For FL under requirements of privacy preservation and straggler mitigation, to overcome the drawbacks of existing CFL methods as mentioned above, we propose a new method named adaptive coded federated learning (ACFL). In ACFL, before training, each device uploads a local coded dataset, which is a transformed version of its local dataset with added random noise. Upon receiving local coded datasets from the devices, the central server generates the global coded dataset. Then, during each training iteration, each non-straggler device computes local gradient based on the current global model and its local dataset, subsequently transmitting this gradient to the central server. After receiving local gradients from the non-straggler devices, the central server estimates the global gradient by aggregating the received gradients and the gradient computed from the global coded dataset, applying aggregation weights determined by an adaptive policy. Under this policy, the performance in terms of privacy and learning is optimized, where the learning performance is described by convergence analysis and the privacy performance is measured by mutual information differential privacy (MI-DP). Finally, in order to verify the superiority of the proposed method, simulations results are provided, where we compare the performance of ACFL and the non-adaptive methods.
@@
    We analytically analyze the trade off between learning and privacy performance of the proposed method.
@@
    Through simulations, we demonstrate that the proposed method surpasses the non-adaptive methods, which attains better learning performance while maintaining the same level of privacy concerns.
@@ II Problem Model
where $\mathbf{W}\in\mathbb{R}^{d\times o}$ is the model parameter matrix, and $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm. Here, we focus on linear regression problems under the FL setting for two reasons. First, linear regression is a very important machine learning model that has attracted significant attention recently due to its ability to make scientific and reliable predictions. As a statistical procedure established a long time ago, its properties are well understood, enabling the rapid training of linear regression models in different applications [26, 27, 28, 29]. Second, nonlinear problems can be transformed into linear regression problems through kernel embedding methods, demonstrating the broad applicability of linear regression models, as indicated by [30].
@@
For the problem introduced above under the FL setting, the training process involves multiple iterations. In each iteration, each device computes local gradient based on its local dataset and transmits it to the central server. After aggregating the received gradients, the central server updates the global model and broadcasts the model parameters to all devices. Over the iterations, due to various incidents, some devices may become stragglers, which are temporarily unresponsive, fail to implement computations, and do not send any messages to the central server, unlike the normally functioning devices [13]. It is assumed that the probability of each device being a straggler is $p$ in each iteration, and the straggler behavior of the devices is independent across iterations and among different devices [19]. It is obvious that the presence of stragglers may degrade the learning performance or even leads to very poor convergence.
@@
To deal with the stragglers in FL, the CFL paradigm has been studied. In CFL, the local dataset $\left({\mathbf{X}^{(i)},\mathbf{Y}^{(i)}}\right)$ is encoded into $\left({\mathbf{H}_{X}^{(i)},\mathbf{H}_{Y}^{(i)}}\right)$ and transmitted to the central server by device $i$ before training begins, $\mathbf{H}_{X}^{(i)}\in\mathbb{R}^{n_{X}^{(i)}\times d}$, $\mathbf{H}_{Y}^{(i)}\in\mathbb{R}^{n_{Y}^{(i)}\times o}$, $\forall i$. The central server utilizes the local coded datasets to generate a global coded dataset, denoted as $\left(\mathbf{\tilde{H}}_{X},\mathbf{\tilde{H}}_{Y}\right)$, instead of retaining all the received coded datasets to save its storage [24]. Subsequently, the training process begins, involving multiple iterations. In each iteration, each non-straggler device computes the gradient based on its local dataset and transmits it to the central server. Upon receiving these messages from the non-stragglers, the central server updates the current model parameter matrix using both the received gradients and the gradient computed from the global coded dataset [24]. At the end of each iteration, the central server broadcasts the updated model parameters to all devices. In CFL, the core issue is to design the method of generating global coded dataset before training and aggregating gradients at the central server during the training process, so that we can achieve a better performance in terms of privacy and learning. A recently proposed method within the CFL framework is known as SCFL [24], which generates local coded datasets by adding noise and random projections. The global dataset is then assembled as the sum of these local coded datasets. Throughout each training iteration, the central server employs a weighted sum with fixed and pre-specified weights across all iterations to aggregate the gradients. However, in SCFL, consistently using fixed and pre-specified weights for aggregating gradients without considering the generation process of the global coded dataset and the dynamic nature of the trained model over iterations results in an unnecessary performance loss.
To overcome the limitations of the existing work, our aim is to attain a better performance in terms of privacy and learning for FL problems under the requirements of privacy preservation and straggler mitigation.
@@ III-A The first stage: before training
Before the training starts, each device uploads a coded dataset to the central server, which is a common practice under the framework of CFL. In the literature, methods commonly used to generate local coded dataset include random matrix projection and noise injection [21, 24, 25]. Random matrix projection reduces the volume of data transmission by lowering the dimension of the projected data compared to the raw dataset. Meanwhile, noise injection enables privacy enhancement at adjustable levels. Inspired by this, in ACFL, the devices encode the local datasets by adding noise to a transformed version of the local dataset. To be more specific, device $i$ uploads coded dataset $\left({\mathbf{H}_{X}^{(i)},\mathbf{H}_{Y}^{(i)}}\right)$ given as
@@
to the central server, where ${\mathbf{N}}_{1}^{\left(i\right)}\in{\mathbb{R}^{d\times d}}$ and ${\mathbf{N}}_{2}^{\left(i\right)}\in{\mathbb{R}^{d\times o}}$ denote the additive noise. In ${\mathbf{N}}_{1}^{\left(i\right)}$, all the elements are independent and identically distributed (i.i.d.) random variables drawn from Gaussian distribution $\mathcal{N}\left({0,\sigma_{1}^{2}}\right)$. Similarly, all the elements in ${\mathbf{N}}_{2}^{\left(i\right)}$ are i.i.d. random variables following distribution $\mathcal{N}\left({0,\sigma_{2}^{2}}\right)$. Differing from existing methods that adopt random matrix projection [21, 24, 25], ACFL projects the local dataset $\mathbf{X}^{(i)}$ using its own transpose. This approach still reduces the volume of transmitted data compared to sending the raw dataset, since $d<m_{i}$ always holds. Furthermore, as will be detailed in Section III-B, it can be noted that $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ and $\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}$ contain information equivalent to that in the raw datasets for computing gradients in linear regression problems. This implies the efficiency of projecting the local dataset as described in (2) and (3).

### v2

@@ Abstract
In this article, we address the problem of federated learning in the presence of stragglers. For this problem, a coded federated learning framework has been proposed, where the central server aggregates gradients received from the non-stragglers and gradient computed from a privacy-preservation global coded dataset to mitigate the negative impact of the stragglers. However, when aggregating these gradients, fixed weights are consistently applied across iterations, neglecting the generation of the global coded dataset and the dynamic nature of the trained model over iterations. This oversight may result in diminished learning performance. To overcome this drawback, we propose a new method named adaptive coded federated learning (ACFL). In ACFL, before the training, each device uploads a local coded dataset with additive noise to the central server to generate a global coded dataset under privacy-preservation requirements. During each iteration of the training, the central server aggregates the gradients received from the non-stragglers and the gradient computed from the global coded dataset, where an adaptive policy for varying the aggregation weights is designed. Under this policy, we optimize the performance in terms of privacy and learning, where the learning performance is analyzed through convergence analysis and the privacy performance in sharing local coded datasets with the server is characterized via mutual information differential privacy. Finally, we perform simulations to demonstrate the superiority of ACFL compared with the baseline methods.
@@ I Introduction
In intelligent systems, edge devices are generating an increasing amount of data, which can be utilized to train machine learning models for various tasks across a wide range of applications, such as smart cities [1], autonomous driving [2], and intelligent healthcare [3]. Traditionally, these data are uploaded to a central server for centralized machine learning. However, uploading datasets directly to the central server without privacy-preservation measures raises severe privacy concerns for the edge devices [4, 5]. As an alternative to centralized machine learning, federated learning (FL) has emerged as a highly effective tool [6, 7, 8, 9]. Under the FL framework, during each iteration of the training process, local devices compute gradients with their local datasets and send these gradients to a central server. The central server then aggregates the received gradients to update the global model [10]. In this manner, the edge devices only share locally computed gradients with the central server, significantly enhancing the privacy of raw datasets.
@@
In FL, the training process may be hindered by devices known as stragglers, which are significantly slower than others or completely unresponsive due to various incidents [11, 12, 13]. Additionally, in FL, the local datasets on different devices can be highly heterogeneous, meaning they are not independent and identically distributed (non-i.i.d.) across devices [14, 15, 16, 17]. In this case, the presence of stragglers can bias the trained global model due to the absence of the contributions of some devices in each iteration. In other words, stragglers can negatively impact the learning performance of FL, potentially leading to poor convergence. In traditional distributed learning scenarios, the challenges posed by stragglers can be addressed through well-known gradient coding (GC) strategies [18, 19, 20, 21, 22, 23, 24]. The core concept of GC is replicating the training data samples and distributing them redundantly across the devices prior to training. During each training iteration, non-straggler devices encode the locally computed gradients and transmit the encoded vectors to the central server. Upon receiving these vectors, the central server can reconstruct the exact global gradient by leveraging the redundancy in the training data distribution. In this way, the impact of stragglers is mitigated, as the missing information from stragglers can be compensated for using the encoded vectors transmitted by non-stragglers. However, in FL scenarios, privacy concerns make it impractical to adopt GC techniques and to allocate training data to devices redundantly in a carefully planned manner. This is because devices are reluctant to share their local datasets directly with others to create data redundancy considering that their local datasets may contain sensitive or private information [25]. Another line of research focused on mitigating stragglers in FL is the development of asynchronous FL algorithms, where the server updates the global model without waiting for all devices to report in every iteration. Instead, the server incorporates stale gradients received from slower devices, thereby decoupling the update process from synchronous coordination. For instance, in [26], a buffered asynchronous aggregation method is proposed, where the server maintains a buffer and performs a model update only when a predefined number of messages from local devices have been collected. In [27], a periodic aggregation protocol is introduced that jointly considers training data representation and channel qualities. However, the performance gain of asynchronous methods is inherently limited, since the information from stragglers is less likely to be frequently utilized by the server, and the missing or stale information from stragglers negatively impacts overall learning performance.
@@
For FL problems, to evade the negative impact of stragglers while avoiding the drawbacks of the aforementioned schemes, a new paradigm named coded federated learning (CFL) has been investigated [28, 29, 30, 31]. In CFL, each device uploads a coded version of its dataset to the central server to form a global coded dataset before training begins. During the training process, the central server aggregates gradients received from the non-stragglers and gradient computed from the global coded dataset to mitigate the negative impact of the stragglers. Compared with sharing raw datasets as in GC, by carefully designing the way of encoding local datasets in CFL, privacy concerns can be largely reduced, and one can even achieve perfect privacy. In addition, compared with asynchronous FL methods, the learning performance can be enhanced in CFL based on the fact that the missing information from the stragglers during the training iterations can be compensated by the information gained from the global coded dataset at the server. The key issue in CFL is how to generate the global coded dataset before training, as well as how to aggregate the gradients at the central server during the training process, in order to achieve better performance in terms of learning and privacy.
@@
There has been some recent progress under the framework of CFL. To be more specific, CFL was firstly investigated in [28] for linear regression problems under the FL setting, where local coded datasets are constructed at the devices by applying random linear projections. In [29], the CFL scheme was combined with distributed kernel embedding to make CFL applicable for nonlinear FL problems. Later, to further enhance the privacy performance of CFL in sharing local coded datasets with the server, a differentially private CFL method was proposed by adding noise to the local coded datasets, where the differential privacy guarantee was proved [30]. Very recently, a stochastic coded federated learning (SCFL) method has been proposed [31]. In SCFL, before the training starts, the devices generate and upload privacy-preservation coded datasets to the central server by adding noise to the projected local datasets, which are used by the central server to generate the global coded dataset. In each iteration of the training process, the central server aggregates gradients received from the non-stragglers and gradient computed from the global coded dataset with fixed aggregation weights. Compared with previous methods, the advantages of SCFL lie in ensuring that the global gradient applied by the central server to update the global model is an unbiased estimation of the true gradient, and providing the trade-off between the privacy performance and the learning performance analytically. However, in SCFL, when the central server aggregates the gradients, it consistently adopts pre-specified and fixed weights across iterations, which neglects the generation process of the global coded dataset and the dynamic nature of the trained model over iterations. This oversight results in an unnecessary performance loss, preventing SCFL from achieving the optimal performance in terms of privacy and learning.
Figure 1: The trade-off between the learning and privacy performance under non-adaptive aggregation weights and adaptive aggregation weights, where the non-adaptive aggregation weight of the gradient computed by the server based on the global coded dataset is fixed at 0.1.
@@
For FL under requirements of privacy preservation and straggler mitigation, to overcome the drawbacks of existing CFL methods as mentioned above, we propose a new method named adaptive coded federated learning (ACFL). In ACFL, before training, each device uploads a local coded dataset, which is a transformed version of its local dataset with additive random noise. Upon receiving local coded datasets from the devices, the central server generates the global coded dataset. Then, during each training iteration, each non-straggler device computes local gradient based on the current global model and its local dataset, subsequently transmitting this gradient to the central server. After receiving local gradients from the non-straggler devices, the central server estimates the global gradient by aggregating the received gradients and the gradient computed from the global coded dataset, applying aggregation weights determined by an adaptive policy. Under this policy, the performance in terms of privacy and learning is optimized, where the learning performance is described by convergence analysis and the privacy performance in sharing local coded datasets with the server is measured by mutual information differential privacy (MI-DP). As an illustration, Fig. 1 presents the relationship between learning performance and privacy performance based on our theoretical analysis, comparing adaptive aggregation weights with non-adaptive (i.e., fixed) weights. The $x$-axis represents privacy performance, where a smaller value indicates higher privacy level. The $y$-axis represents learning performance, where a smaller value indicates faster convergence (See Fig. 4 in Section IV-C for more descriptions.). It can be observed that by using adaptive aggregation weights determined by the adaptive policy, better learning performance is achieved for the same level of privacy performance, compared to the case of using fixed weights, as done in existing methods such as SCFL. Finally, in order to verify the superiority of the proposed method, simulations results are provided, where we compare the performance of ACFL and the baseline methods.
It is worth noting that although both ACFL and SCFL adopt the CFL framework, there are significant differences between them, which are outlined below:
  1. 1.
    ACFL and SCFL employ different methods for encoding local datasets at the devices and generating the global coded dataset at the server. Unlike SCFL, which uses a separate projection matrix to transform local datasets for encoding, ACFL projects each local dataset using its own transpose. The advantage of this approach is that the transformed local datasets retain information equivalent to the raw datasets for gradient computation, leading to better learning performance at the same privacy level. This rationale will be explained in more detail in Section III and further verified in Section V.
  2. 2.
    Unlike SCFL, which consistently uses pre-specified and fixed weights across iterations, ACFL fully considers the generation process of the global coded dataset and the dynamic nature of the trained model over iterations and uses adaptive aggregation weights to aggregate gradients at the server for model updates. The use of adaptive aggregation weights ensures optimal performance in terms of both privacy and learning for the proposed method.
  3. 3.
    Due to the very different implementations, the theoretical analysis presented in this paper is novel and distinct from that in SCFL, even though both methods use the same measure, MI-DP, to characterize the privacy level of the FL system in sharing local coded datasets with the server.
@@
    We analyze the trade-off between learning and privacy performance of the proposed method in sharing local coded datasets with the server.
@@
    Through simulations, we demonstrate that the proposed method surpasses the baseline methods, which attains better learning performance under certain privacy levels.
@@ II Problem Model
where $\mathbf{W}\in\mathbb{R}^{d\times o}$ is the model parameter matrix, and $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm. Here, we focus on linear regression problems under the FL setting for two reasons. First, linear regression is a very important machine learning model that has attracted significant attention recently due to its ability to make scientific and reliable predictions. As a statistical procedure established a long time ago, its properties are well understood, enabling the rapid training of linear regression models in different applications [33, 34, 35, 36]. For instance, in agricultural applications, linear regression models can be used to predict rice cultivation timing based on weather data. In medical applications, simulated scenarios of traditional Chinese bone-setting techniques provide valuable data for analysis. By applying linear regression, one can establish relationships between measurable features of the manipulation process and the underlying conceptual knowledge, enabling the extraction of abstract insights from observable physical actions [33]. Second, nonlinear problems can be transformed into linear regression problems through techniques such as high-order approximations or distributed kernel embedding [29], demonstrating the broad applicability of linear regression models. In addition, many classification problems can be reformulated as linear regression problems as well. For instance, binary classification tasks can be approached using linear regression by encoding class labels numerically, e.g., 0 and 1, and interpreting the regression output as a probability estimate.
@@
For the problem introduced above under the FL setting, the training process involves multiple iterations. In each iteration, the devices compute local gradients based on their local datasets and transmit to the central server. After aggregating the received gradients, the central server updates the global model and broadcasts the model parameters to all devices. Over the iterations, due to various incidents, some devices may become stragglers, which are temporarily unresponsive, fail to implement computations, and do not send any messages to the central server, unlike the normally functioning devices [18]. It is assumed that the probability of each device being a straggler is $p$ in each iteration, and the straggler behavior of the devices is independent across iterations and among different devices [24]. This assumption corresponds to practical scenarios in which devices are equipped with hardware or resources of similar quality, and stragglers arise due to temporary computation or communication issues, such as transient hardware errors [37]. These types of errors are typically independent across devices and over time, and occur with roughly equal likelihood on each device [37]. Additionally, the probability of becoming a straggler is assumed to remain constant, which is reasonable when the behavior of devices is relatively stationary during the training process [^1] The methodology and analysis presented in this paper can be extended to scenarios in which straggler behavior is heterogeneous or correlated across devices, as well as to cases where the straggler behavior is time-varying or exhibits temporal correlation across iterations due to resource dependency. However, such extensions are beyond the primary focus of this work, which are left for future extensions. [18].
The presence of stragglers may degrade learning performance or even lead to poor convergence. As an illustrative example, Fig. 2 shows the model update process with and without stragglers. When there are no stragglers, the global model is consistently updated in the direction of steepest descent. In contrast, in the presence of stragglers, missing gradients from the stragglers introduce gradient misalignment, which destabilizes the optimization trajectory and slows convergence. To be more specific, as shown in Fig. 2, after the same number of iterations starting from the same initial model, the optimal model can be reached in the absence of stragglers, where the training loss is minimized. However, in the presence of stragglers, the resulting model remains far from the optimum, indicating that only a suboptimal solution is obtained. In this case, additional iterations are required to reach the optimal model, compared to the scenario without stragglers.
Figure 2: An illustrative example showing the effects of the stragglers.
@@
To deal with the stragglers in FL, the CFL paradigm can be employed. In CFL, the local dataset $\left({\mathbf{X}^{(i)},\mathbf{Y}^{(i)}}\right)$ is encoded into $\left({\mathbf{H}_{X}^{(i)},\mathbf{H}_{Y}^{(i)}}\right)$ and transmitted to the central server by device $i$ before training begins, $\mathbf{H}_{X}^{(i)}\in\mathbb{R}^{n_{X}^{(i)}\times d}$, $\mathbf{H}_{Y}^{(i)}\in\mathbb{R}^{n_{Y}^{(i)}\times o}$, $\forall i$. The central server utilizes the local coded datasets to generate a global coded dataset, denoted as $\left(\mathbf{\tilde{H}}_{X},\mathbf{\tilde{H}}_{Y}\right)$, instead of retaining all the received coded datasets to save its storage [31]. Subsequently, the training process begins, involving multiple iterations. In each iteration, each non-straggler device computes the gradient based on its local dataset and transmits it to the central server. Upon receiving these messages from the non-stragglers, the central server updates the current model parameter matrix using both the received gradients and the gradient computed from the global coded dataset [31]. At the end of each iteration, the central server broadcasts the updated model parameters to all devices. In CFL, the core issue is to design the method of generating global coded dataset before training and aggregating gradients at the central server during the training process, so that we can achieve a better performance in terms of privacy and learning. A recently proposed method within the CFL framework is known as SCFL [31]. However, in SCFL, consistently using fixed and pre-specified weights for aggregating gradients without considering the generation process of the global coded dataset and the dynamic nature of the trained model over iterations results in an unnecessary performance loss. Based on this, we aim to overcome the limitations of existing work and achieve better performance in terms of both privacy and learning, while fully exploiting the advantages of the CFL paradigm.
@@ III-A The first stage: before training
Before the training starts, each device uploads a coded dataset to the central server, which is a common practice under the framework of CFL. In the literature, methods commonly used to generate local coded datasets include matrix projection and noise injection [28, 31, 32]. Matrix projection reduces the volume of data transmission by lowering the dimension of the projected data compared to the raw datasets. Meanwhile, noise injection enables privacy enhancement at adjustable levels. Inspired by this, in ACFL, the devices encode the local datasets by adding noise to a transformed version of the local datasets. To be more specific, device $i$ uploads coded dataset $\left({\mathbf{H}_{X}^{(i)},\mathbf{H}_{Y}^{(i)}}\right)$ given as
@@
to the central server, where ${\mathbf{N}}_{1}^{\left(i\right)}\in{\mathbb{R}^{d\times d}}$ and ${\mathbf{N}}_{2}^{\left(i\right)}\in{\mathbb{R}^{d\times o}}$ denote the additive noise. In ${\mathbf{N}}_{1}^{\left(i\right)}$, all the elements are independent and identically distributed (i.i.d.) random variables drawn from Gaussian distribution $\mathcal{N}\left({0,\sigma_{1}^{2}}\right)$. Similarly, all the elements in ${\mathbf{N}}_{2}^{\left(i\right)}$ are i.i.d. random variables following distribution $\mathcal{N}\left({0,\sigma_{2}^{2}}\right)$. Differing from existing methods that adopt random matrix projection [28, 31, 32], ACFL projects each local dataset using its own transpose. This approach still reduces the volume of transmitted data compared to sending the raw datasets, since $d<m_{i}$ always holds. Furthermore, as will be detailed in Section III-B, $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ and $\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}$ contain information equivalent to that in the raw datasets for computing gradients in linear regression problems. This implies the efficiency of projecting the local datasets as described in (2) and (3). Furthermore, it is worth emphasizing that the computation of $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ and $\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}$ during the encoding operation on the devices incurs almost no additional computational load. This is because, in the absence of the encoding operation, $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ and $\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}$ would still need to be computed later by each device to obtain the local gradients during the second stage of training iterations, as described in Section III-B.